OpenStudy (anonymous):

An aircraft, used for fire spotting, flies from its base to locate a fire at an unknown distance, x km away. It travels straight to the fire and back, averaging 240 km/h for the outward trip and 320 km/h for the return trip. If the plane was away for 35 minutes, find the distance, x km.

OpenStudy (alivejeremy):

we have 2 elements here 1)departure element s= v=x/t(1) 2)arrival elements = v`=x`/t`(2) v=240 km/h v`=320 km/h t+t`=35 min = 35/60 hour x=x` ( it means the distance is the same ) as we named the formulas (1) and (2) see below (1)+(2) = v+v` = x/t+x/t` = xt`+xt / t.t` = x( t+t`) / t.t` ====> x . 35/60 / t.t` = 560 ==> x / t.t` = 960 ==> v / t` = 960 ==> t` = v / 960==> t`=0.25 you have t` and also v` right now so you can calculate the x v`=x/t` ==> x = t`v`==> x = 320 * 0.25 = 80 km

OpenStudy (alivejeremy):

Yw :)

OpenStudy (alivejeremy):

@jenbubbly

OpenStudy (anonymous):

THANK YOU so much @alivejeremy

OpenStudy (alivejeremy):

Yw :)))